3. Let L be the length of the curve x=e^y from (1,0) to (e,1), integrating with respect to x an integral expression for L is_____.

4. Let be the length of the curve y=ln(x) from (1,0) to (e,1), integrating with respect to y, an integral expression for L is_____.
1
Expert's answer
2019-12-09T12:42:54-0500

3.

Changing the parameter of the curve to "x" we get the expression

"y = \ln x,\; x \in [1,e]."

Applying the formula of function graph arc length we get

"L = \int_1^e \sqrt{1 + (\ln'(x))^2} dx = \int_1^e \sqrt{1 + \frac{1}{x^2}} dx \\=
\int_1^e \frac{\sqrt{1+x^2}}{x} dx."


Answer: "L = \int_1^e \frac{\sqrt{1+x^2}}{x} dx."


4.


Similarly, let us change the curve parameter to "y"

"x = e^y,\;y \in [0,1]."

Then,

"L = \int_0^1 \sqrt{1 + (\frac{d}{dy}e^y)^2}dy = \int_0^1 \sqrt{1 + e^{2y}}dy."


Answer: "L = \int_0^1 \sqrt{1 + e^{2y}}dy."